Wesley Taylor is considering investing some money that he inherited. The following payoff table gives the profits that would be realized during the next year for each of three investment alternatives Wesley is considering:


STATE OF Nature
DECISION
ALTERNATIVE

GOOD

ECONOMY

POOR

ECONOMY


Stock Market

$110,000

-50,000


Bonds

   40,000    30,000

CDs
   43,000 43,000
Probability 0.6 0.4


REQUIRED:
a.   What decision would maximize expected profit?
b.   What is the maximum amount that should be paid for a perfect forecast of the economy?
Answer #1

Ans a)

The EMV or Expected monetary value of 'Stock market' decision alternative = 0.6 * 110,000 + 0.4 * (-50,000) = $46,000

The EMV or Expected monetary value of 'Bonds' decision alternative = 0.6 * 40,000 + 0.4 * 30,000 = $36,000

The EMV or Expected monetary value of 'CDs' decision alternative = 0.6 * 43,000 + 0.4 * 43,000 = $43,000

The maximum expected profit is from Stock market decision. Hence Wesley Taylor should invest in stock market.

Ans b)

The EVPI or expected value of perfect information = Expected value with perfect information - Expected value without perfect information

The Expected value without perfect information is the maximum expected monetary value out of the given alternatives. In this case, it is 46,000.

EVPI = 0.6 * 110,000 + 0.4 * 43,000 - 46,000

EVPI = $83,200 - $46,000

EVPI = $37,200

Hence the maximum amount that should be paid for a perfect forecast of the economy is $37,200.